calculate the density of hydrogen bromide (HBr) gas in grams per liter at 733 mmHg and 46°C



1
Expert's answer
2022-01-10T13:56:23-0500

P=pressure of gas in atm = 733 / 760 atm = 0.9644 atm

T = Temperaturein kelvin = ( 273 + 46 ) k = 319 k

Molar mass of HBr = M = 81 g / mol .


PM = dRT .


d = PM / RT .

= 0.9644 × 81 / 0.0821 × 319

= 2.9826 gram / litre.